the speed of an automobile in mph varies directly with its speed in kph. a speed of 64 mph is equivalent to a speed of 103 kph. write a linear equation that relates speed in miles per hour to speed in kph
Answer by stanbon(75887)

the speed of an automobile in mph varies directly with its speed in kph. a speed of 64 mph is equivalent to a speed of 103 kph. write a linear equation that relates speed in miles per hour to speed in kph
----------------
mph = k*kph
---
Solve for "k":
64 = k*103
k = 64/103
----
mph = (64/103)kph
==========================
Cheers,
Stan H.
============